What is the smallest perimeter that can be used to make a rectangle with an area of 256cm squared?
Answer by stanbon(75887)

What is the smallest perimeter that can be used to make a rectangle with an area of 256cm squared?
-----------------------------
Area = LW
256 = LW
W = (256/L)
-------
Perimeter = 2(L+W)
P = 2(L + (256/L))
---
P = 2(L^2+256)/L
----
Take the derivative to get:
P' = 2[(L^2+256)-L(2L)]/L^2
----
Solve: 2[(L^2+256)-L(2L)]/L^2 = 0
---
Solve: L^2+256-2L^2 = 0
-L^2 = -256
L^2 = 256
---
L = sqrt(256) = 16
Therefore W = sqrt(256) = 16
===============================
Perimeter = 4*16 = 64
===============================
Cheers,
Stan H.